how do i find KWH from WATTs, let's say i have 381.6933966 [w] device, how do I find the cost per KWH if the rate is $.12?
Answer #1

Answer :- Watt (W) is the unit of power and kilo-watt-hour (kWh) is the unit for energy. 1 W power for 1 Hr of time consumes 1 Whr of energy or 0.001 kWh of energy.

We have power value as 381.6933966 W, for per KWh i.e. for 1 Hr of time, the energy consumed is-

E = 381.6933966 / 1000 kWh = 0.3816933966 kWh.

So the total cost per kWh is = $0.12 * 0.3816933966 = $0.0458.